# Sequential Attention-based Sampling for Histopathological Analysis

### v2

@@ Abstract
Deep neural networks are increasingly applied for automated histopathology. Yet, whole-slide images (WSIs) are often acquired at gigapixel sizes, rendering it computationally infeasible to analyze them entirely at high resolution. Diagnostic labels are largely available only at the slide-level, because expert annotation of images at a finer (patch) level is both laborious and expensive. Moreover, regions with diagnostic information typically occupy only a small fraction of the WSI, making it inefficient to examine the entire slide at full resolution. Here, we propose SASHA – Sequential Attention-based Sampling for Histopathological Analysis – a deep reinforcement learning approach for efficient analysis of histopathological images. First, SASHA learns informative features with a lightweight hierarchical, attention-based multiple instance learning (MIL) model. Second, SASHA samples intelligently and zooms selectively into a small fraction (10-20%) of high-resolution patches, to achieve reliable diagnosis. We show that SASHA matches state-of-the-art methods that analyze the WSI fully at high-resolution, albeit at a fraction of their computational and memory costs. In addition, it significantly outperforms competing, sparse sampling methods. We propose SASHA as an intelligent sampling model for medical imaging challenges that involve automated diagnosis with exceptionally large images containing sparsely informative features.
@@ 1 Introduction
Deep learning models have emerged as a major frontier in automated medical imaging diagnosis. Such models have been successfully applied to identify tumorous tissue in whole-slide histopathological images (WSI) [13, 9, 11, 18]. Yet, WSIs are often acquired at gigapixel sizes, and hardware and memory bottlenecks render it impractical to analyze these images directly. Moreover, analyzing WSI-s at high resolution is inefficient because regions with diagnostic information (e.g., tumor cells) make up only a small fraction of the entire image, and are not clearly identifiable at low resolutions. As a result, methods that predict slide-level labels from low-resolution WSI images suffer from performance bottlenecks [26, 15]. By contrast, methods that incorporate finer-grained “patch-level” information perform better [25, 33]. Yet, these latter methods require significant resources, needing expert annotations of WSI images at the patch level.
The Multiple Instance Learning (MIL) framework [4], addresses these limitations elegantly. In MIL, the entire WSI is divided into non-overlapping patches, with each patch being encoded into a high-dimensional feature vector. The feature vectors are then aggregated using simple methods, like max- or mean-pooling, and this “bag” of features is labeled at the slide level. Furthermore, diagnostic regions in the image can be identified with weighted aggregation methods, like Attention-Based deep Multiple Instance Learning (ABMIL) [14], which assigns attention scores to each patch based on their putative relevance, prior to aggregation.
Figure 1: SASHA: An attention-augmented deep RL model. Starting with low-resolution WSI features $S_{0}\in\mathbb{R}^{N\times d}$, at each time step $t$, the deep RL agent selects a patch $a_{t}$ (red outline) to zoom in at high-resolution, based on policy $\pi(a_{t}|S_{t})$. High-resolution features $V(a_{t})\in\mathbb{R}^{d}$ are extracted and aggregated using a Heirarchical Attention-based Feature Distiller (HAFED, Section 3.2). The states $S_{t}$ of patches with features similar to $a_{t}$ are updated with a Targeted State Updater (TSU, Section 3.3). At the episode end, the classifier (Section 3.2) predicts the presence or type of cancer.
Despite these advancements, a major drawback of MIL approaches is the need to analyze a significant fraction of the WSI patches at high resolution. Because tumor patches are sparsely distributed in the WSI, feature extraction, especially for the vast majority of non-tumor patches, can be expensive and increases inference time as well as GPU memory requirements. To address this challenge, Zhao et al [36] proposed a deep reinforcement learning (RL)-based approach – RLogist – that examines the WSI at low resolution and zooms in to specific high-resolution patches whose feature representations can be selectively analyzed to predict labels at the slide-level.
@@
Here, we address these limitations by combining RL with multi-attention MIL in a framework we call SASHA - Sequential Attention-based Sampling for Histopathological Analysis (Fig. 1). Our study makes the following key contributions:
@@
We show that SASHA achieves state-of-the-art performance on two cancer benchmarks [3, 23], comparable to models that analyze the entire WSI at high resolution, while requiring only $\sim 6\%$ of their memory for WSI representation and with upto 8$\times$ faster inference times.
@@ 2 Related Work
Deep learning approaches for automated histopathology have a rich and detailed history (reviewed in [31, 24]). Here, we review specifically those methods that inspire key aspects of our solution and, therefore, form benchmarks for performance comparisons.
@@
Multiple Instance Learning (MIL). MIL is a weakly supervised learning paradigm widely applied in WSI analysis. Unlike fully supervised learning, where each instance is associated with a label, MIL involves dividing the slide into a “bag” of instances (patches); the model is then trained with bag-level labels. MIL is, therefore, suitable for classification problems where slide-level, but not patch-level, labels are available. Early MIL approaches relied on instance selection to identify the most relevant patches within a bag ([8]). Recent models, such as attention-based MIL (ABMIL) [14], assign importance weights to patches, allowing the model to focus on diagnostically relevant regions.
@@
Advances in attention-based MIL. A key limitation of these approaches is that each patch is treated independently of the others, even though many may share strong feature correlations. Recent models exploit correlations among instance representations using a self-attention mechanism (e.g., TransMIL [29]). Other approaches cluster instances based on their attention scores [20]. Yet, these attention mechanisms may capture only some types of diagnostic features, leading to overfitting and lack of generalization. To overcome this, Attention-Challenging Multiple Instance Learning (ACMIL) [35] employs multiple attention blocks, each focusing on distinctive diagnostic features. Inspired by these approaches, our model incorporates multiple attention heads for diagnostically-relevant feature extraction and a correlation-based state update rule for efficient feature updates.
@@
Deep Reinforcement Learning (DRL) in medical imaging. Conventional RL learns a policy for actions that maximizes the expected reward in novel, unstructured environments. DRL augments RL with the representational power of deep neural networks. The power of DRL was originally demonstrated in challenging strategy and video games (e.g., Deep-Q, AlphaGo), [22, 30]. DRL is particularly useful in domains with limited annotations, such as automated medical diagnosis, and has been successfully applied to various medical imaging tasks, including localization of anatomical landmarks in medical scans [1], lesion detection [21, 32], 3D image segmentation of MRI scans [19] and tumor classification in WSIs [36, 37]. In such tasks, a DRL agent learns an optimal search strategy by exploring different portions of the environment (here, image) and receiving sparse feedback in the form of classification accuracy, without the need for analyzing the entire image at high resolution.
@@
DRL for histopathology. Diagnosis of WSIs by human histopathologists involves viewing the slide – at a low (“scanning”) resolution – and zooming into suspected tumor regions – at high resolution – for reliable diagnosis. In this context, a recent approach, RLogist [36], models the scanning process of histopathologists as a Markov Decision Process (MDP) using an RL agent. The agent learns an efficient observation strategy, selectively zooming into only a few high-resolution WSI patches before making a classification decision, thereby reducing computational overhead and inference time. Similarly, PAMIL [37] addresses the shortcomings of instance-based clustering methods using an RL agent to form pseudo-bags (aggregates) of related instances from high-resolution WSIs. It models pseudo-bag formation as an MDP, forming new pseudo-bags based on past knowledge from the previously formed bags, informed by the label error arising from bag-level predictions.
@@
Feature extraction for histopathology. Deep neural networks such as ResNet [12] and Vision Transformers (ViT) [6], pre-trained on ImageNet [5], are commonly used for feature extraction in computer vision tasks. Their exposure to millions of natural images enables them to identify generic features that encode the image parsimoniously, for specialized downstream tasks. Yet, pretraining on natural images is not ideal for medical images, due to differences in their feature distributions. [2]. For instance, histopathology images lack a canonical orientation, exhibit low color variation due to standardized staining protocols, and their diagnostic information varies with magnification. To address this, Kang et al. [16] introduced ResNet-50 and ViT-S models with self-supervised pretraining on large WSI datasets (e.g., TCGA, TULIP).
@@ 3.1 Problem Statement and Overview
For our task, the model seeks to predict binary target labels $Y\in\{0,1\}$ for a given Whole Slide Image (WSI) $X$. The latter is a 3-channel RGB image that is available in the dataset at multiple different levels of magnification, whereas the former either indicates the presence (versus absence) of cancer or indicates one of two subtypes. We notate the size (in pixels) of the image at magnification level $m\in\{1,2,3\}$ as $G^{m}$. The goal of our approach is to scan the image at the lowest resolution $G^{3}$, which we call the “scanning-level” as in [36], while zooming into and analyzing only select portions of the image at a higher resolution $G^{1}$. For this, each WSI at $G^{3}$ is divided into multiple non-overlapping patches, such that a given WSI $X$ is represented as $X=\{x_{1},\ldots,x_{N}\}$, where $N$ is the total number of patches at low resolution $G^{3}$; note that $N$ can vary across different WSI images. Now, when zoomed in, these patches are magnified to a high-resolution; for this, we choose the highest resolution $G^{1}$ in all datasets tested. Therefore, each patch $x_{i}$ at the low-resolution comprises of $k$ patches at the high-resolution; we denote these as $\{x_{i1},\ldots,x_{ik}\}$.
@@
Given these definitions, the three main components of the algorithm are as follows (Fig. 1):
@@
Hierarchical Feature Distiller: This hierarchical attention-based network extracts diagnostic features from the $k$ high-resolution patches for each low-resolution patch and creates a $d$ dimensional feature representation for it. The output of this network is used to update WSI “state” and to train a downstream WSI label classifier; see Section 3.2 for details.
@@
Targeted State Updater: This network maintains a state of the entire WSI, comprising a $d$ dimensional feature representation for each low-resolution patch. As each patch is zoomed in to high-resolution, the state vector is updated, but only for instances whose features are correlated with those of the sampled patch; see Section 3.3 for details.
@@
Deep RL Agent: This agent comprises a policy network and a value network. Based on the current state, the former network produces a probability distribution over a set of actions from which the agent selects the next patch to zoom in, whereas the latter network is used to compute a generalized advantage estimate for each state-action combination. Both networks are optimized using the Proximal Policy Optimization (PPO) algorithm; see Section 3.4 for details.
@@
Preprocessing and Feature Extraction. Prior to training and testing, we preprocess the WSI and extract features using a pre-trained model, as follows: To isolate histopathological tissue from the glass slide background on the WSI we apply established segmentation methods [20] at low resolution. The extracted tissue is then divided into non-overlapping patches of size $(256\times 256\times 3)$, again at both resolutions; these operations were performed using the code provided as part of the CLAM [20] GitHub repository. For the first level of feature extraction, we utilize a pre-trained ViT-based encoder [6], as described in the Related Work (Section 2). Each input patch is encoded into a feature embedding vector of dimension ${d}$. We denote the input feature representation of a WSI at low-resolution as $Z\in\mathbb{R}^{N\times d}$, and at high-resolution as $U\in\mathbb{R}^{N\times k\times d}$.
@@ 3.2 Hierarchical Attention-based Feature Distiller (HAFED)
A key first step with diagnosis of large WSIs is to learn informative, yet parsimonious, feature representations for the high-resolution patches. Such representations are critical for the RL agent to update WSI state, as it makes decisions about which patch to visit next, as well as for the final classification. Recent advances in attention-based MIL models compute such informative representations for high-resolution whole-slide images (WSIs) [14]. However, in an RL setting, the state representation for the WSI at each time step comprises a mix of information both at low and high-resolutions, depending on the specific patches zoomed in (see next). It is, therefore, desirable for both the low-resolution and high-resolution feature representations to have matched dimensionality; such a design choice considerably simplifies downstream computations (e.g., classification). To achieve this, we propose a Hierarchical Attention-based FEature Distiller (HAFED) (Fig. 2a).
HAFED is a two-stage attention-based model designed to operate directly on high-resolution feature patches while producing representations compatible with low-resolution patches. In the first stage (Fig. 2a, Feature Aggregator), the model takes as input high-resolution features $U\in\mathbb{R}^{N\times k\times d}$, and learns to estimate attention scores for each of the $k$ sub-patches, reflecting the latter’s relative importance in each of the $N$ low resolution patches; a stochastic instance masking strategy is applied to avoid over-fitting [35]. A weighted aggregation of the sub-patches based on these attention scores produces a compressed feature representation $V\in\mathbb{R}^{N\times d}$, thereby reducing the dimensionality of the high-resolution features and aligning them with the dimensionality of the low-resolution features $Z\in\mathbb{R}^{N\times d}$. In the second stage (Fig. 2a, Classifier), a second attention mechanism operates across the $N$ patch-level features in $V$ to identify diagnostically important patches at the slide level. A weighted aggregation of the $N$ features, guided by the second-stage attention scores, yields a final slide-level embedding $h\ \in\mathbb{R}^{d}$ that summarizes the entire WSI. This $d$-dimensional embedding is then used for downstream WSI classification. We employed multiple ($M$) attention branches, and the model is trained with the sum of three losses [35]: i) a similarity loss to learn distinct attention weight patterns in each branch, ii) the label loss associated with each branch, and iii) overall label loss.
During training, both stages of the model are trained end-to-end by visiting all of the $N$ patches at high resolution. Thus, attention weights, and feature representations, in both the first and second stage of the model are learned in a “label-aware” manner using the ACMIL loss (classifier and dissimilarity losses). During inference, a single high-resolution patch $a_{t}$ is visited at each timestep $t$ and its compressed feature representation $V(a_{t})\in\mathbb{R}^{d}$ is used to update the slide state. While previous approaches [36] employ generic, label-agnostic feature extractors HAFED identifies important regions at both the sub-patch (first-stage) and patch (second-stage) levels, which enables efficient learning of diagnostically-relevant and parsimonious feature representations from high-resolution WSIs.

### v3

@@ Abstract
Deep neural networks are increasingly applied in automated histopathology. Yet, whole-slide images (WSIs) are often acquired at gigapixel sizes, rendering them computationally infeasible to analyze entirely at high resolution. Diagnostic labels are largely available only at the slide-level, because expert annotation of images at a finer (patch) level is both laborious and expensive. Moreover, regions with diagnostic information typically occupy only a small fraction of the WSI, making it inefficient to examine the entire slide at full resolution. Here, we propose SASHA – Sequential Attention-based Sampling for Histopathological Analysis – a deep reinforcement learning approach for efficient analysis of histopathological images. First, SASHA learns informative features with a lightweight hierarchical, attention-based multiple instance learning (MIL) model. Second, SASHA samples intelligently and zooms selectively into a small fraction (10-20%) of high-resolution patches to achieve reliable diagnoses. We show that SASHA matches state-of-the-art methods that analyze the WSI fully at high resolution, albeit at a fraction of their computational and memory costs. In addition, it significantly outperforms competing, sparse sampling methods. We propose SASHA as an intelligent sampling model for medical imaging challenges that involve automated diagnosis with exceptionally large images containing sparsely informative features [^1] Model implementation is available at: https://github.com/coglabiisc/SASHA.
@@ 1 Introduction
Deep learning models have emerged as a major frontier in automated medical imaging diagnosis. For example, such models accurately identify tumorous tissue in whole-slide histopathological images (WSI) [hou2016patch, gecer2018dcnc, hameed2020breast, li2021multi]. Yet, hardware and memory bottlenecks render it impractical to analyze WSIs acquired at gigapixel sizes. Moreover, analyzing WSI-s at high resolution is inefficient because regions with diagnostic information (e.g., tumor cells) make up only a small fraction of the entire image, and are not clearly identifiable at low resolutions. As a result, methods that predict slide-level labels from low-resolution WSIs suffer from performance bottlenecks [ruan2021btl1, jenkinson2024btl2]. By contrast, methods that incorporate fine-grained “patch-level” information perform better [roy2019patch, wang2021tumor]. Yet, these latter methods require significant resources and expert annotations of WSIs at the patch level.
The Multiple Instance Learning (MIL) framework [das2018multiple], addresses these limitations elegantly. In MIL, the WSI is divided into non-overlapping patches, with each patch being encoded into a high-dimensional feature vector. The feature vectors are aggregated using simple methods, like max- or mean-pooling, and this “bag” of features is slide-level labels. Furthermore, diagnostic regions can be identified with weighted aggregation methods, like Attention-Based deep Multiple Instance Learning (ABMIL) [ilse2018abmil], which assigns patch-level attention scores, based on their relevance.
Figure 1: SASHA: An attention-augmented deep RL model. Starting with low-resolution WSI features $S_{0}\in\mathbb{R}^{N\times d}$, at each time step $t$, the deep RL agent selects a patch $a_{t}$ (red outline) to zoom in at high-resolution, based on policy $\pi(a_{t}|S_{t})$. High-resolution features $V(a_{t})\in\mathbb{R}^{d}$ are extracted and aggregated using a Heirarchical Attention-based Feature Distiller (HAFED, Section˜3.2). The states $S_{t}$ of patches with features similar to $a_{t}$ are updated with a Targeted State Updater (TSU, Section˜3.3). At the end of each episode, the classifier (Section˜3.2) predicts the presence or type of cancer.
Nonetheless, a major limitation of MIL-based approaches is the need to process most WSI patches at high resolution. Because tumors typically occupy only a small portion of the slide, feature extraction for the majority of non-tumor patches becomes computationally inefficient, increasing inference time and memory usage. To address this, Zhao et al. [zhao2022rlogist] proposed a deep reinforcement learning (RL)-based approach – RLogist, that first examines WSIs at low resolution and then selectively zooms into diagnostically relevant high-resolution patches to predict slide-level labels.
@@
Here, we address these limitations by combining RL with multi-attention MIL in a framework we call SASHA - Sequential Attention-based Sampling for Histopathological Analysis (Fig.˜1). Our study makes the following key contributions:
@@
[…]
@@ 2 Related Work
Deep learning approaches for automated histopathology have a rich and detailed history (reviewed in [srinidhi2021deep, Qu_2022_DL_histo]). Here, we review specifically those methods that inspire key aspects of our solution and, therefore, form benchmarks for performance comparisons.
@@
Multiple Instance Learning (MIL). MIL is a weakly supervised learning paradigm widely applied in WSI analysis. Unlike fully supervised learning, where each instance is associated with a label, MIL involves dividing the slide into a “bag” of instances (patches); the model is then trained with bag-level labels. MIL is, therefore, suitable for classification problems where slide-level, but not patch-level, labels are available. Early MIL approaches relied on instance selection to identify the most relevant patches within a bag ([GADERMAYR2024_MIL_review]). Recent models, such as attention-based MIL (ABMIL) [ilse2018abmil], assign importance weights to patches, allowing the model to focus on diagnostically relevant regions.
@@
Advances in attention-based MIL. A key limitation of these approaches is that each patch is treated independently of the others, even though many may share strong feature correlations. Recent models exploit correlations among instance representations using a self-attention mechanism (e.g., TransMIL [shao2021transmil]). Other approaches cluster instances based on their attention scores [lu2021clam]. Yet, these attention mechanisms may capture only some types of diagnostic features, leading to overfitting and lack of generalization. To overcome this, Attention-Challenging Multiple Instance Learning (ACMIL) [zhang2023acmil] employs multiple attention blocks, each focusing on distinctive diagnostic features. Inspired by these approaches, our model incorporates multiple attention heads for diagnostically relevant feature extraction and a correlation-based state update rule for efficient feature updates. Moreover, earlier approaches largely process every patch at high resolution, which is computationally inefficient. ZoomMIL [zoom_mil] addresses this by recursively sampling a fixed number of patches per resolution based on top-$k$ attention scores. However, its sampling strategy depends on predefined attention criteria, which motivates the use of adaptive and context-aware patch selection methods.
@@
Deep Reinforcement Learning (DRL) in medical imaging. Conventional RL learns a policy for actions that maximizes the expected reward in novel, unstructured environments. DRL augments RL with the representational power of deep neural networks. The power of DRL was originally demonstrated in challenging strategy and video games (e.g., Deep-Q, AlphaGo), [mnih2013playingatariDRL, silver2017mastering]. DRL is particularly useful in domains with limited annotations, such as automated medical diagnosis, and has been successfully applied to various medical imaging tasks, including localization of anatomical landmarks in medical scans [ALANSARY2019156], lesion detection [Maicas2017lesion, stember2020lesion], 3D image segmentation of MRI scans [liao2019segmentation] and tumor classification in WSIs [zhao2022rlogist, zheng2024cvpr]. In such tasks, a DRL agent learns an optimal search strategy by exploring different portions of the environment (here, image) and receiving sparse feedback in the form of classification accuracy, without the need for analyzing the entire image at high resolution.
@@
DRL for histopathology. Diagnosis of WSIs by human histopathologists involves viewing the slide – at a low (“scanning”) resolution – and zooming into suspected tumor regions – at high resolution – for reliable diagnosis. In this context, a recent approach, RLogist [zhao2022rlogist], models the scanning process of histopathologists as a Markov Decision Process (MDP) using an RL agent. The agent learns an efficient observation strategy, selectively zooming into only a few high-resolution WSI patches before making a classification decision, thereby reducing computational overhead and inference time. Similarly, PAMIL [zheng2024cvpr] addresses the shortcomings of instance-based clustering methods using an RL agent to form pseudo-bags (aggregates) of related instances from high-resolution WSIs. It models pseudo-bag formation as an MDP, forming new pseudo-bags based on past knowledge from the previously formed bags, informed by the label error arising from bag-level predictions.
@@
Feature extraction for histopathology. Deep neural networks such as ResNet [he2016deep] and Vision Transformers (ViT) [doso2020vit], pre-trained on ImageNet [imagenet], are commonly used for feature extraction in computer vision tasks. Their exposure to millions of natural images enables them to identify generic, relevant features, for specialized downstream tasks. Yet, pretraining on natural images is not ideal for medical image diagnosis. [REMEDIS]. For instance, histopathology images lack a canonical orientation, exhibit low color variation due to standardized staining protocols, and their diagnostic information varies with magnification. To address this, Kang et al. [kang2023ssl] introduced ResNet-50 and ViT-S models with self-supervised pretraining on large WSI datasets (e.g., TCGA, TULIP).
@@ 3.1 Problem Statement and Overview
For our task, the model seeks to predict binary target labels $Y\in\{0,1\}$ for a given Whole Slide Image (WSI) $X$. The latter is a 3-channel RGB image that is available in the dataset at multiple different levels of magnification, whereas the former either indicates the presence (versus absence) of cancer or indicates one of two subtypes. We notate the size (in pixels) of the image at magnification level $m\in\{1,2,3\}$ as $G^{m}$. The goal of our approach is to scan the image at the lowest resolution $G^{3}$, which we call the “scanning-level” as in [zhao2022rlogist], while zooming into and analyzing only select portions of the image at a higher resolution $G^{1}$. For this, each WSI at $G^{3}$ is divided into multiple non-overlapping patches, such that a given WSI $X$ is represented as $X=\{x_{1},\ldots,x_{N}\}$, where $N$ is the total number of patches at low resolution $G^{3}$; note that $N$ can vary across different WSI images. Now, when zoomed in, these patches are magnified to a high-resolution; for this, we choose the highest resolution $G^{1}$ in all datasets tested. Therefore, each patch $x_{i}$ at the low-resolution comprises of $k$ patches at the high-resolution; we denote these as $\{x_{i1},\ldots,x_{ik}\}$.
@@
Given these definitions, the three main components of the algorithm are as follows (Fig.˜1):
@@
Hierarchical Feature Distiller: This hierarchical attention-based network extracts diagnostic features from the $k$ high-resolution patches for each low-resolution patch and creates a $d$ dimensional feature representation for it. The output of this network is used to update WSI “state” and to train a downstream WSI label classifier; see Section˜3.2 for details.
@@
Targeted State Updater: This network maintains a state of the entire WSI, comprising a $d$ dimensional feature representation for each low-resolution patch. As each patch is zoomed in to high-resolution, the state vector is updated, but only for instances whose features are correlated with those of the sampled patch; see Section˜3.3 for details.
@@
Deep RL Agent: This agent comprises a policy network and a value network. Based on the current state, the former network produces a probability distribution over a set of actions from which the agent selects the next patch to zoom in, whereas the latter network is used to compute a generalized advantage estimate for each state-action combination. Both networks are optimized using the Proximal Policy Optimization (PPO) algorithm; see Section˜3.4 for details.
@@
Preprocessing and Feature Extraction. Prior to training and testing, we preprocess the WSI and extract features using a pre-trained model, as follows: To isolate histopathological tissue from the glass slide background on the WSI we apply established segmentation methods [lu2021clam] at low resolution. The extracted tissue is then divided into non-overlapping patches of size $(256\times 256\times 3)$, again at both resolutions; these operations were performed using the code provided as part of the CLAM [lu2021clam] GitHub repository. For the first level of feature extraction, we utilize a pre-trained ViT-based encoder [doso2020vit], as described in the Related Work (Section˜2). Each input patch is encoded into a feature embedding vector of dimension d{d}. We denote the input feature representation of a WSI at low-resolution as $Z\in\mathbb{R}^{N\times d}$, and at high-resolution as $U\in\mathbb{R}^{N\times k\times d}$.
@@ 3.2 Hierarchical Attention-based Feature Distiller (HAFED)
A key first step with diagnosis of large WSIs is to learn informative, yet parsimonious, feature representations for the high-resolution patches. Such representations are critical for the RL agent to update WSI state, as it makes decisions about which patch to visit next, as well as for the final classification. Recent advances in attention-based MIL models compute such informative representations for high-resolution whole-slide images (WSIs) [ilse2018abmil]. However, in an RL setting, the state representation for the WSI at each time step comprises a mix of information both at low and high-resolutions, depending on the specific patches zoomed in (see next). It is, therefore, desirable for both the low-resolution and high-resolution feature representations to have matched dimensionality; such a design choice considerably simplifies downstream computations (e.g., classification). To achieve this, we propose a Hierarchical Attention-based FEature Distiller (HAFED) (Fig.˜2a).
[…]
During training, both stages of the model are trained end-to-end by visiting all of the $N$ patches at high resolution. Thus, attention weights, and feature representations, in both the first and second stage of the model are learned in a “label-aware” manner using the ACMIL loss (classifier and dissimilarity losses). During inference, a single high-resolution patch $a_{t}$ is visited at each timestep $t$ and its compressed feature representation $V(a_{t})\in\mathbb{R}^{d}$ is used to update the slide state. While previous approaches [zhao2022rlogist] employ generic, label-agnostic feature extractors HAFED identifies important regions at both the sub-patch (first-stage) and patch (second-stage) levels, which enables efficient learning of diagnostically-relevant and parsimonious feature representations from high-resolution WSIs.